OpenStudy (driftracer305):

solve 2 log 3 x = 54

OpenStudy (anonymous):

is it \[2\log_3(x)=54\]?

OpenStudy (driftracer305):

yea........ it is base 3

OpenStudy (driftracer305):

log 3(x) = 27.......?

OpenStudy (anonymous):

thought so divide both sides by two and start with \[\log_3(x)=27\] then it should be fairly obvious

OpenStudy (driftracer305):

x = 9.....? my only doubt?

OpenStudy (anonymous):

on no

OpenStudy (driftracer305):

ohh wait...

OpenStudy (driftracer305):

it is 3^27........?

OpenStudy (anonymous):

\[\log_b(x)=y\iff b^y=x\]they are trying to trick you

OpenStudy (anonymous):

yes, the answer is \(x=3^{27}\)

OpenStudy (anonymous):

no don't fall for the trick!

OpenStudy (driftracer305):

but how come the answer is 7.6256.......... whereas 3^27 is a huge #

OpenStudy (anonymous):

probably written in scientific notation let me check

OpenStudy (driftracer305):

ok

OpenStudy (abmon98):

1) Divide both sides by 2 \[\log_{3} x=54/2\]\[\log_{3} x=27\]\[a=\log_{b} (b^a)\]\[\log_{3} x=\log_{3} (3^(27)\] x=3^27

OpenStudy (anonymous):

how'd i guess? \[3^{27}=7.625597484987 × 10^{12}\]

OpenStudy (driftracer305):

yea but @Abmon98 ........by that we get 7625597........ and so on..... but the answer here is only 7.6256

OpenStudy (anonymous):

scientific notation is being used

OpenStudy (driftracer305):

@satellite73 ............ hmm i see........ they didnt ask for scientific notation...... and the answer only states 7.6256........... that's it..... not even raised to something..... so i was wondering....??

OpenStudy (anonymous):

i don't know what you mean by "the answer states"

OpenStudy (driftracer305):

ohh @satellite73 ....... i meant that answer only says 7.6256 and nothing like 7.6 x 10^12...............but dont worry....i got it.........bdw thx a lot for all the help..!!

OpenStudy (triciaal):

I don't have the solution i've forgotten a lot but I know that when you multiply numbers you add the logs. 54 can be expressed as 2 *27 which is 2* 3^3 so it would be Log base 3 of 2 plus log base 3 of 3^3 3 log to base 3 of 3 = 3